# Transmon qubit modeling and characterization for Dark Matter search

### v2

@@ II Circuit design and simulation
where $p_{i}$ and $\tan{\delta_{i}}$ are the participation ratios of the material and geometry-dependent loss tangents, respectively. Estimating the participation ratio in thin layers, such as the oxide regrowth between material interfaces, requires accurate modeling of the electric field near metal edges. This is challenging to achieve with three-dimensional simulations because they require extremely fine-grained meshing in the region of interest, often making such simulations too inefficient for standard desktop computing. This problem can be addressed using analytical models [43, 44] or a hybrid 3D-2D simulation strategy [45].
@@ III Experimental characterization
Table 2: Qubit characterization results, showing the agreement between measured and expected values.
@@
The coherence times of both qubits fall significantly below expectations. The experimental values for $T_{1}$, $T_{2}$, and $T_{2}^{*}$, reported in Table 2 are the average results of several individual measurements taken over around $50$ hours. The expected values for $T_{1}$ account for dielectric losses and Purcell decay through the resonator. The quality factors considered for this estimate include the dielectric qubit quality factor $Q_{\text{TLS}}$ estimated in Sec. II and a resonator quality factor estimated from measured internal and coupling quality factors $Q_{r}=(1/Q_{i}+1/Q_{c})^{-1}$.

### v3

@@ I Introduction
Our proposal addresses the challenge of introducing a magnetic field into the experimental setup, similar to the approach described in [18], while incorporating modifications to enable quantum non-demolition (QND) detection. In the method presented in [18], photons produced in the haloscope via the axion conversion process exit the haloscope and propagate toward a photon counter. This system employs a frequency-tunable cavity that efficiently absorbs resonant photons, leading to qubit excitation via a four-wave mixing process enabled by a dedicated pump line. In contrast to this detection mechanism, which is destructive, we propose replacing the aforementioned photon counter with a QND-compatible detector, in analogy with the one developed in [19]. Although the setup in [19] employs a three-dimensional (3D) storage cavity, in our work two-dimensional (2D) cavities are explored. This alternative would offer a more compact, fully on-chip implementation, despite reducing the quality factor. Notably, the proposed device shares the same architecture as that used in the direct excitation method; however, certain lines are repurposed, as the four-wave mixing process is no longer required. The system consists of a transmon qubit coupled to a readout resonator, a high-quality tunable storage cavity, and a dedicated driveline. Photons originating from the haloscope and traveling along the transmission line are trapped within the storage cavity when resonant, with a lifetime $\tau=Q_{s}/\omega_{s}$, where $Q_{s}$ is the storage cavity quality factor and $\omega_{s}$ the resonant frequency. During this time, the qubit state undergoes precession according to Eq. 2. Provided that $\tau$ is significantly longer than the qubit readout time, multiple non-destructive measurements can be carried out, achieving dark count suppression. Fig. 1 illustrates the proposed detection scheme.
Figure 1: Schematic representation of a QND axion detection scheme. The axion field $a$, interacting with the magnetic field $\vec{B}$ inside the haloscope, produces a photon ($\gamma$) and a virtual photon ($\gamma^{*}$). The photon is absorbed by a tunable storage cavity (green), inducing a phase precession in the qubit (blue) state. Repeated photon detection is achieved by performing parity measurements via the readout resonator (red).
@@ II Circuit design and simulation
where $p_{i}$ and $\tan{\delta_{i}}$ are the participation ratios of the material and geometry-dependent loss tangents, respectively. These ratios are determined by the fraction of energy stored within a volume enclosing the Josephson element relative to the total energy stored in the qubit-resonator system. Estimating the participation ratio in thin layers, such as the oxide regrowth between material interfaces, requires accurate modeling of the electric field near metal edges. This is challenging to achieve with three-dimensional simulations because they require extremely fine-grained meshing in the region of interest, often making such simulations too inefficient for standard desktop computing. This problem can be addressed using analytical models [43, 44] or a hybrid 3D-2D simulation strategy [45].
@@ III Experimental characterization
Table 2: Qubit characterization results, showing the agreement between measured and expected values. $T^{*}_{2}$ and $T_{2}$ refer to Ramsey and Hahn-echo characterization experiments [38], respectively.
@@
[…]